integrate the following integral
∫ x sqrt(x+2) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below


I = \int x\sqrt{x+2}dx
I = \int (x+2-2)\sqrt{x+2}dx
I = \int (x+2)^{3/2}dx-2\int \sqrt{x+2}dx
I = \frac{2(x+2)^{5/2}}{5}-2\int \sqrt{x+2}dx
I = \frac{2(x+2)^{5/2}}{5}-\frac{4(x+2)^{3/2}}{3}+c